# Theory of enhanced-by-coincidence neural information transmission

### v1

@@ B.1 Single, inhomogenenous Poisson process with Ornstein-Uhlenbeck rate
We consider an inhomogeneous Poisson process whose rate is $r(t)=\mu+\chi(t)$. The fluctuations around its average are such that $\chi(t)\right>_{\chi}=0$, and $\chi(t)\chi(t^{\prime})\right>_{\chi}$ is a given function of $|t-t^{\prime}|$, and where $_{\chi}$ is the average over realisations of the noise $\chi$. Let $\rho(t)=\sum_{k}\delta(t-t_{k})$ be the density of spikes of such process, where $t_{k}$ is the $k$-th spike. In other words, $\rho(t)=\dot{n}(t)$ where $n(t)$ is the number of spikes $t_{k}<t$ in a single realisation (conditioned to a fixed realisation of $r$), and the dot stands for the temporal derivative. We are interested in the two-time correlation function, or the covariance
@@ B.2 Many encoding neurons
We now consider a multivariate Poisson process whose ${N}-$dimensional vector of rates ${\bf x}(t)={\bm{\mu}}+{\bm{\chi}}(t)$ whose fluctuations around its mean $\bm{\mu}$ satisfy the Ornstein-Uhlenbeck process: $\dot{\bm{\chi}}(t)=-{\bm{\chi}}(t)/{\tau_{\rm c}}+(2C/{\tau_{\rm c}})^{1/2}\cdot{\bm{\eta}}(t)$, where 1/2 is the matrix square root, and where $C$ is a covariance matrix. Indeed, the stationary distribution for the solutions $\bm{\chi}$ of this Ornstein-Uhlenbeck process is a multivariate Gaussian with null mean and covariance equal to $C$. We define as well the Pearson correlation matrix $\bar{C}$ as: $\bar{C}_{ij}=C_{ij}(v_{i}v_{j})^{-1/2}$, where the $v_{i}:=C_{ii}$’s are the stationary variances of the $\chi_{i}(t)$’s. Finally, we define the spike count variance of the $j$-th neuron, $\sigma_{j}^{2}:=2{\tau_{\rm c}}v_{j}$. Following the same reasoning lines of the previous subsection, we learn that, if $\rho_{i}(t)$ is the firing rate of the $i$-th Poisson process, it is:
@@
${\sigma_{\rm c}}$ $={\bf w}^{\dagger}\cdot C\cdot{\bf w}.$ (24d)
@@ B.3 The Valente et al 2021 model
B.4 A difference between the present notations and that of Valente et al
@@ Appendix C Numerical methods
In the main text, we have presented a stochastic-processes theory of the Valente et al model of coincidence-detection defined in section II. The diffusion approximation (see III.1) allows us to treat the membrane potential of the readout LIF neuron as the solution of a stochastic differential equation. In the article, we compare the approximated analytical solutions to such equations, with numerical simulations, that help us to asses the approximations’ validity limit. We have, in particular, and as in references [17, 19], numerically solved Eqs. (5) using the Euler-Maruyama method [60, 61], completed with a numerical protocol (see below) that corrects the bias induced by the finiteness of the Number of Integration Steps (NIS) per unit time, ${n_{\rm is}}$.
In brief, our numerical estimation for the rate mean and variance, ${r_{\rm R}},{v_{\rm R}}$ respectively, are performed as follows. First, we employ the Euler-Maruyama method with ${T_{\rm c}}{n_{\rm is}}$ integration steps to numerically integrate Eqs. (5) with the desired set of parameters ${\mu_{\rm I}},{\sigma_{\rm I}},{\sigma_{\rm c}},{\tau_{\rm c}},{\tau_{\rm m}}$ in a time interval of length ${T_{\rm c}}$. We perform a number of ${n_{\rm tr}}$ different realisations of such integration, differing in the independent random samples of the white noises $\eta,\eta^{\prime}$, and in the initial conditions for $\xi(0),v(0)$ (that can either be independently sampled from their stationary distributions, or chosen, as for our numerical results, to coincide with those of the previous realisation in the last instant value). For each realisation, we store the corresponding sequence of deadout neuron spike times $\vec{t}$. Afterwards, we compute the across-trials average and variance of the rate ${n_{\rm R}}({T_{\rm c}})/{T_{\rm c}}$, where ${n_{\rm R}}({T_{\rm c}})=\text{len}(\vec{t})$ in single realisations, and where ${T_{\rm c}}=20$. These are, respectively, our finite-${n_{\rm is}}$ estimations for the rate mean and variance, $\tilde{r_{\rm R}},\tilde{v_{\rm R}}$ (circles and squares in Figs. 4,7,8-A,8-B,8-C). Importantly, and as already specified in the main article text, all the rate variance estimations that we report are rate variances per unit time: ${v_{\rm R}}{T_{\rm c}}$. Indeed, both the singular and the regular part of the temporal correlation function $\sf C$ induce rate variances that decreases as $1/{T_{\rm c}}$: to eliminate this dependency, we report the variance in intervals of unit length. Consequenly, the SNR analised in Secs. IV.1,IV.2 is the SNR per unit time window, or: ${T_{\rm c}}^{-1/2}\times\text{SNR}$.
@@ C.2 Error estimation of rate mean, variance, and SNR
As said in the main text, the statistical uncertainty (error-bars) of our estimation $\tilde{r_{\rm R}}$ in Figs. 4,8-A is computed as the standard deviation of the rate ${n_{\rm R}}({T_{\rm c}})/{T_{\rm c}}$ across the ${n_{\rm tr}}$ trials (hence, multiplied by ${T_{\rm c}}^{1/2}$). The errors of our estimations of the rate variance in Figs. 7,8-B are the confidence intervals of the Chi-squared distribution with ${n_{\rm tr}}=10^{3}$ degrees of freedom and a probability of $\alpha_{\chi}=0.95$. The errors of the SNR in Fig. 8 correspond to the standard deviation across bootstrap realisations of the ${n_{\rm tr}}$ realisations of spike times –for each of which we compute the SNR (see [48]). Finally, the uncertainties of the NIS-corrected rates (and $\beta_{r}$, see below) are the errors of the regression (see [48] for details).
@@ C.3 Parameter fixing
We have performed two sets of numerical simulations: the 1st and the 2nd sets of simulations. The 1st set of simulations is to assess the numerical dependence of the SNR with ${\mu_{\rm I}},{\tau_{\rm c}},{v_{\rm c}}({\alpha_{\rm V}}),{n_{\rm is}}$. It is the set that we use to perform the NIS-corrected estimations of the SNR as a function of ${\tau_{\rm c}}$ and ${\alpha_{\rm V}}$ (in Figs. 4-A,6,7,8-A,8-B,8). The 2nd set of simulations is for a fixed value of ${n_{\rm is}}$ and ${\alpha_{\rm V}}$, and varying ${\mu_{\rm I}},{\alpha_{\rm V}}$. We use it to assess the differences between the Moreno et al, and quenched-noise approximations in Fig. 4-A (use use, indeed, a larger value of ${v_{\rm V}}$, to have larger values of $\alpha_{\rm M}={\sigma_{\rm c}}^{2}/{\sigma_{\rm I}}^{2}$ given ${\tau_{\rm c}}$ and, hence, to explore the region in which Moreno solution is expected to be less accurate than the quenched-noise approximation).
In tables 3,4 we report the paramter values of each set (please, see [48] as well).
In both sets, given ${\alpha_{\rm V}},{\tau_{\rm c}}$ and ${v_{\rm V}}$, we consequently initialize ${v_{\rm c}}$ through Eq. (7b) and, consequently, ${\sigma_{\rm c}}^{2}=2{\tau_{\rm c}}{v_{\rm c}}$. Analogously, given $w$ and ${\mu_{\rm I}}$, we initialize ${\sigma_{\rm I}}=(w{\mu_{\rm I}})^{1/2}$. The number of encoding neurons ${N}=20$ is fixed for both simulations, and plays the only role of changing the relation between ${v_{\rm c}}$ and ${\alpha_{\rm V}}$ in Eq. (7b).
Finally, for Figs. 2,3 we have performed shorter, ad-hoc, single-realisation simulations. There, we simulate the encoding spike trains as well, for visualisation purposes only. In Figs. 2,3, we first integrate a solution ${\bf x}(t)=(x_{1}(t),x_{2}(t))$ of a two-dimensional Ornstein-Uhlenbeck process for the the encoding rate; we therefore generate the encoding spike trains from an inhomogeneous Poisson distribution with corresponding inhomogeneous rate ${\bf x}(t)$ (for visualisation only, not to give them as an input to the LIF membrane potential $V(t)$); we finally integrate the LIF membrane potential $V(t)$ according to Eq. (5), using the rate ${\bf x}(t)$ (times a constant, $w$) at the place of ${\mu_{\rm I}}+{\sigma_{\rm c}}\xi(t)$ in Eq. (5), consistently with the diffusion approximation.
@@ C.4 Algorithm of finite-NIS effect reduction
Eventual completions of this work could include a comparison with alternative numerical schemes of integration of stochastic differential equations beyond Euler-Maruyama, as those based on the path integral [58, 59]).
@@ C.5 Numerical estimation of the temporal distribution function $g$
In Fig. 12 we illustrate the emerging $g$’s for two values of ${\mu_{\rm I}}$ and of ${\tau_{\rm c}}$. We can see how, indeed, for large enough values of $t$, $g(t)$ becomes statistically compatible with $\tilde{r_{\rm R}}$. This cannot be achieved without correcting for the finite-${T_{\rm c}}$ bias.
Figure 12: Numerical estimation of the readout neuron temporal distribution function $g(t)$. Different curves correspond to different values of ${\mu_{\rm I}}$ and of ${\tau_{\rm c}}$ (see the figure key) in the 1st simulation set, with ${\alpha_{\rm V}}=0.9$ and ${n_{\rm is}}=2^{18}$. The horizontal lines are the NIS-corrected estimations for ${r_{\rm R}}$ corresponding to each ${\mu_{\rm I}},{\tau_{\rm c}}$.
@@ Appendix D Validity range of the approximations
The membrane potential of the LIF readout neuron in the Valente model is described, in the diffusion approximation, by the system of equations (5). Such approximation consists in approximating the statistics of the number of encoding spikes in the time window $\Delta$ (sampled from a Poisson distribution with average $\Lambda=\int_{\Delta}{\rm d}t\,{\bf w}\cdot{\bm{\rho}}(t)$) by the statistics of the normal distribution. For the diffusion approximation to be correct, the number of encoding spikes in relevant time windows should be $\gg 1$. In Fig. 13, we show ${\mu_{\rm I}}/{r_{\rm R}}$, which is the average number of encoding spikes in windows of length equal to the average readout ISI, ${r_{\rm R}}^{-1}$, for all the probed values of ${\mu_{\rm I}}$ and ${\tau_{\rm c}}$, and for both values of ${\alpha_{\rm V}}$. We observe that the average number of encoding spikes in the average readout neuron ISI is, in the worst case, larger than $100$, and larger than $1000$ in the better case. This is consistent with the diffusion approximation.

### v2

@@ B.1 Single, inhomogenenous Poisson process with Ornstein-Uhlenbeck rate
We consider an inhomogeneous Poisson process whose rate is $r(t)=\mu+\chi(t)$. The fluctuations around its average are such that $\chi(t)\right>_{\chi}=0$, and $\chi(t)\chi(t^{\prime})\right>_{\chi}$ is a given function of $|t-t^{\prime}|$, and where $_{\chi}$ is the average over realizations of the noise $\chi$. Let $\rho(t)=\sum_{k}\delta(t-t_{k})$ be the density of spikes of such process, where $t_{k}$ is the $k$-th spike. In other words, $\rho(t)=\dot{n}(t)$ where $n(t)$ is the number of spikes $t_{k}<t$ in a single realization (conditioned to a fixed realization of $r$), and the dot stands for the temporal derivative. We are interested in the two-time correlation function, or the covariance
@@ B.2 Many encoding neurons
We now consider a multivariate Poisson process whose ${N}-$dimensional vector of rates ${\bf x}(t)={\bm{\mu}}+{\bm{\chi}}(t)$ is such that the fluctuations around the mean $\bm{\mu}$ satisfy the Ornstein-Uhlenbeck process: $\dot{\bm{\chi}}(t)=-{\bm{\chi}}(t)/{\tau_{\rm c}}+(2C/{\tau_{\rm c}})^{1/2}\cdot{\bm{\eta}}(t)$, where 1/2 is the matrix square root, and where $C$ is a covariance matrix. Indeed, the stationary distribution for the solutions $\bm{\chi}$ of this Ornstein-Uhlenbeck process is a multivariate Gaussian with null mean and covariance equal to $C$. We define as well the Pearson correlation matrix $\bar{C}$ as: $\bar{C}_{ij}=C_{ij}(v_{i}v_{j})^{-1/2}$, where the $v_{i}:=C_{ii}$’s are the stationary variances of the $\chi_{i}(t)$’s. Finally, we define the spike count variance of the $j$-th neuron, $\sigma_{j}^{2}:=2{\tau_{\rm c}}v_{j}$. Following the same reasoning lines of the previous subsection, we learn that, if $\rho_{i}(t)$ is the firing rate of the $i$-th Poisson process, it is:
@@
${\sigma_{\rm c}}^{2}$ $={\bf w}^{\dagger}\cdot C\cdot{\bf w}.$ (24d)
@@ B.3 The Valente et al 2021 model
B.4 Analogy with the random walk problem
In the diffusion approximation, valid for ${\mu_{\rm I}}\gg({\sigma_{\rm I}}^{2}+{\sigma_{\rm c}}^{2})^{1/2}$ (see [42, 17], the stimulus-readout probability distribution $\text{prob}({\vec{t}}|s)$ (on which the SNR depends, through ${r_{\rm R}},{v_{\rm R}}$ in Eq. (4)), becomes the distribution of sequences of first passage times of a continuous-time random walker $V(t)$ (see Eq. (5b)), with damp $1/{\tau_{\rm m}}$, constant drift ${\mu_{\rm I}}$, white noise drift ${\sigma_{\rm I}}\eta(t)$, colored noise drift ${\sigma_{\rm c}}\xi(t)$, and the reset mechanism of the LIF neuron (that conditions the stationary distribution of the $\xi$’s [18]). The readout neuron Inter-Spike Interval (ISI) time is the first passage time of such a random walker (see [49, 65], and references therein).
B.5 A difference between the present notations and that of Valente et al
@@ Appendix C Numerical methods
In the main text, we have presented a stochastic-processes theory of the NERM defined in section II. The diffusion approximation (see III.1) allows us to treat the membrane potential of the readout LIF neuron as the solution of a stochastic differential equation. In the article, we compare the approximated analytical solutions to such equations, with numerical simulations, that help us to asses the approximations’ validity limit. We have, in particular, and as in references [17, 19], numerically solved Eqs. (5) using the Euler-Maruyama method [74, 75], completed with a numerical protocol (see below) that corrects the bias induced by the finiteness of the Number of Integration Steps (NIS) per unit time, ${n_{\rm is}}$.
In brief, our numerical estimation for the rate mean and variance, ${r_{\rm R}},{v_{\rm R}}$ respectively, are performed as follows. First, we employ the Euler-Maruyama method with ${T_{\rm c}}{n_{\rm is}}$ integration steps to numerically integrate Eqs. (5) with the desired set of parameters ${\mu_{\rm I}},{\sigma_{\rm I}},{\sigma_{\rm c}},{\tau_{\rm c}},{\tau_{\rm m}}$ in a time interval of length ${T_{\rm c}}$. We perform a number of ${n_{\rm tr}}$ different realizations of such integration, differing in the independent random samples of the white noises $\eta,\eta^{\prime}$, and in the initial conditions for $\xi(0),v(0)$ (that can either be independently sampled from their stationary distributions, or chosen, as for our numerical results, to coincide with those of the previous realization in the last instant value). For each realization, we store the corresponding sequence of readout neuron spike times $\vec{t}$. Afterwards, we compute the across-trials average and variance of the rate ${n_{\rm R}}({T_{\rm c}})/{T_{\rm c}}$, where ${n_{\rm R}}({T_{\rm c}})=\text{len}(\vec{t})$ in single realizations, and where ${T_{\rm c}}=20$. These are, respectively, our finite-${n_{\rm is}}$ estimations for the rate mean and variance, $\tilde{r_{\rm R}},\tilde{v_{\rm R}}$ (circles and squares in Figs. 5,8,9-A,9-B,9-C). Importantly, and as already specified in the main article text, all the rate variance estimations that we report are rate variances per unit time: ${v_{\rm R}}{T_{\rm c}}$. Indeed, both the singular and the regular part of the temporal correlation function $\sf C$ induce rate variances that decreases as $1/{T_{\rm c}}$: to eliminate this dependency, we report the variance in intervals of unit length. Consequenly, the SNR analised in Secs. IV.2,IV.3 is the SNR per unit time window, or: $\text{SNR}={T_{\rm c}}^{-1/2}\times\text{SNR}({T_{\rm c}})$. The so defined standard deviation of the rate in the unit time interval, $({v_{\rm R}}{T_{\rm c}})^{1/2}$ coincides with the value of the rate standard deviation, ${v_{\rm R}}^{1/2}$ that we obtain if we divide each of our ${n_{\rm tr}}$ realizations of the readout spike trains $\vec{t}_{r}$ in $m={\rm int}({T_{\rm c}})$ chunks, in such a way that we have a new set of $m\times{n_{\rm tr}}$ shorter realizations, of length $1$ each, and we compute the standard deviation of $\text{len}(\vec{t})$ across this novel set of spike trains. The same is true for the SNR per unit time interval.
@@ C.2 Error estimation of rate mean, variance, and SNR
As said in the main text, the statistical uncertainty (error-bars) of our estimation $\tilde{r_{\rm R}}$ in Figs. 5,9-A is computed as the standard deviation of the rate ${n_{\rm R}}({T_{\rm c}})/{T_{\rm c}}$ across the ${n_{\rm tr}}$ trials (hence, multiplied by ${T_{\rm c}}^{1/2}$). The errors of our estimations of the rate variance in Figs. 8,9-B are the confidence intervals of the Chi-squared distribution with ${n_{\rm tr}}=10^{3}$ degrees of freedom and a probability of $\alpha_{\chi}=0.95$. The errors of the SNR in Fig. 9 correspond to the standard deviation across bootstrap realizations of the ${n_{\rm tr}}$ realizations of spike times –for each of which we compute the SNR (see [48]). Finally, the uncertainties of the NIS-corrected rates (and $\beta_{r}$, see below) are the errors of the regression.
@@ C.3 Parameter fixing
We have performed seven sets of numerical simulations. In Table 3 we specify the name, purpose and characteristics of each of such (set of) simulations. Among them, the so called 1st simulation is to assess the dependence of the SNR with ${\mu_{\rm I}},{\tau_{\rm c}},{v_{\rm c}}({\alpha_{\rm V}}),{n_{\rm is}}$. It is the set that we use to perform the NIS-corrected estimations of the SNR as a function of ${\tau_{\rm c}}$ and ${\alpha_{\rm V}}$ (in Figs. 5-A,7,8,9-A,9-B,9). In tables 410 we report the values of the parameters of each set of simulations (see [48] as well). In each set of simulations, given the value of the varying independent parameters, the rest of the parameters are consequently set according to Eqs. (6,7b). Analogously, given $w$ and ${\mu_{\rm I}}$, we set ${\sigma_{\rm I}}=(w{\mu_{\rm I}})^{1/2}$. In the tables, “diffusion” refers to the NERM model in the diffusion approximation, Eq. (5) (for this model, the number of encoding neurons ${N}$ plays the only role of changing the relation between ${v_{\rm c}}$ and ${\alpha_{\rm V}}$ in Eq. (7b)). “diffusion-N” model refers to the NERM in diffusion approximation with the sum of single stochastic differential equations for each encoding neuron as input current (see Sec. D.2). Finally, “spike-enc” refers to the NERM model with spiking encoding input current (see Sec. D.4).
Table 3: Outline of simulations.
simulation name purpose model varying NIS-extrapolation alg. figures
1st SNR estimation in the diffusion NERM vs ${\tau_{\rm c}}$, fixed ${v_{\rm c}}$ diffusion ${\mu_{\rm I}},{\alpha_{\rm V}},{\tau_{\rm c}},{n_{\rm is}}$ yes 5-B,7,8,9,12,13,24,25,26,27,28,29
2nd quenched-noise vs Moreno difference assessment diffusion ${\mu_{\rm I}},{\alpha_{\rm V}}$ no 5-A
3rd SNR estimation in the diffusion NERM vs ${\sigma_{\rm V}}$, fixed ${\tau_{\rm c}}$ diffusion ${\mu_{\rm I}},{\sigma_{\rm V}},{n_{\rm is}}$ no 23
4th Figs. 3,4 SNR estimation diffusion-N ${\mu_{\rm I}},{\alpha_{\rm V}},N$ no 14
5th Figs. 3,4 SNR estimation diffusion ${\mu_{\rm I}},{\alpha_{\rm V}},{\sigma_{\rm V}}$ no 14
6th spike-enc. LIF vs diffusion LIF comparison spike-enc ${\mu_{\rm I}},w$ no 15,16
7th SNR estimation in the spike-enc. NERM vs ${\sigma_{\rm V}}$ spike-enc ${\mu_{\rm I}},{\sigma_{\rm V}}$ no 17,18
@@
${T_{\rm c}}$ no $20$
Table 6: Parameters of the third simulation set.
variable varying value/range
${\mu_{\rm I}}$ yes $1180+10m/$ 11, with $m=0,\ldots,11$
${\sigma_{\rm V}}$ yes $8.9442\times(m/9)$, with $m=0,\ldots,9$
=0. , 1.118025, 2.23605 , 3.354075, 4.4721 , 5.590125, 6.70815 , 7.826175, 8.9442
${n_{\rm is}}$ yes $2^{m}$ for $m=15,\ldots,18$
${\alpha_{\rm V}}$ no $0.9$
${\tau_{\rm c}}$ no $0.1$
${\tau_{\rm m}}$ no $0.005$
$N$ no $20$
$w$ no $0.1$
${n_{\rm tr}}$ no $100$
${T_{\rm c}}$ no $200$
Table 7: Parameters of the fourth simulation set.
variable varying value/range
${\mu_{\rm I}}$ yes $80,90$
${\alpha_{\rm V}}$ yes $0,0.9$
$N$ yes $6,10,20$
${\sigma_{\rm V}}$ yes $={\sigma_{\rm V}}^{(N=20)}\times 20/N$, with ${\sigma_{\rm V}}^{(N=20)}=4.47213$
${n_{\rm is}}$ no $32768$
${\tau_{\rm c}}$ no $0.1$
${\tau_{\rm m}}$ no $0.005$
$w$ no $0.1$
${n_{\rm tr}}$ no $200$
${T_{\rm c}}$ no $200$
Table 8: Parameters of the fifth simulation set.
variable varying value/range
${\mu_{\rm I}}$ yes $80,90$
${\alpha_{\rm V}}$ yes $0,0.9$
${\sigma_{\rm V}}$ yes $0,4.47213$
$N$ no $20$
${n_{\rm is}}$ no $32768$
${\tau_{\rm c}}$ no $0.1$
${\tau_{\rm m}}$ no $0.005$
$w$ no $0.1$
${n_{\rm tr}}$ no $200$
${T_{\rm c}}$ no $200$
Table 9: Parameters of the sixth simulation set.
variable varying value/range
${\mu_{\rm I}}$ yes $1170+130m/$ 11 with $m=0,\ldots,11$
${\sigma_{\rm V}}$ yes $0,4.47213$
$w$ yes $0.05,0.1,0.2$
${\alpha_{\rm V}}$ no $0.9$
$N$ no $20$
${n_{\rm is}}$ no $65536$
${\tau_{\rm c}}$ no $0.1$
${\tau_{\rm m}}$ no $0.005$
${n_{\rm tr}}$ no $20$
${T_{\rm c}}$ no $200$
Table 10: Parameters of the seventh simulation set.
variable varying value/range
${\mu_{\rm I}}$ yes $1540+50m/$ 15 with $m=0,\ldots,15$
${\sigma_{\rm V}}$ yes $0,2,4.47213$
${\alpha_{\rm V}}$ no $0.9$
$N$ no $20$
${n_{\rm is}}$ no $32768$
${\tau_{\rm c}}$ no $0.1$
${\tau_{\rm m}}$ no $0.005$
$w$ yes $0.1$
${n_{\rm tr}}$ no $200$
${T_{\rm c}}$ no $200$
@@ C.4 Algorithm of finite-NIS effect reduction
Future completions of this work could include a comparison with alternative numerical schemes of integration of stochastic differential equations beyond Euler-Maruyama, as those based on the path integral [72, 73].
@@ C.5 Numerical estimation of the temporal distribution function $g$
In Fig. 13 we illustrate the resulting $g$’s for two values of ${\mu_{\rm I}}$ and of ${\tau_{\rm c}}$. We can see how, indeed, for large enough values of $t$, $g(t)$ becomes statistically compatible with $\tilde{r_{\rm R}}$. This cannot be achieved without correcting for the finite-${T_{\rm c}}$ bias.
Figure 13: Numerical estimation of the readout neuron temporal distribution function $g(t)$. Different curves correspond to different values of ${\mu_{\rm I}}$ and of ${\tau_{\rm c}}$ (see the figure key) in the 1st simulation set, with ${\alpha_{\rm V}}=0.9$ and ${n_{\rm is}}=2^{18}$. The horizontal lines are the NIS-corrected estimations for ${r_{\rm R}}$ corresponding to each ${\mu_{\rm I}},{\tau_{\rm c}}$. See the rest of the parameter values in Table 4.
@@ Appendix D Validity range of the approximations
D.1 Validity of the diffusion approximation for the encoding current
The membrane potential of the LIF readout neuron in the Valente model is described, in the diffusion approximation, by the system of equations (5). Such approximation consists in approximating the statistics of the number of encoding spikes in the time window $\Delta$ (sampled from a Poisson distribution with average $\Lambda=\int_{\Delta}{\rm d}t\,{\bf w}\cdot{\bm{\rho}}(t)$) by the statistics of the normal distribution. For the diffusion approximation to the encoding current to be correct, the average encoding current in the diffusion approximation ${\mu_{\rm I}}$, must be large enough in units of the noise amplitude, $({\sigma_{\rm I}}^{2}+{\sigma_{\rm c}}^{2})^{1/2}$ (the square root of the total spike count variance per unit time). If ${\mu_{\rm I}}$ were not large enough in units of $({\sigma_{\rm I}}^{2}+{\sigma_{\rm c}}^{2})^{1/2}$, negative values of the spike count rate could exhibit non-negligible probability and, consequently, one would overestimate the count variance. We have assessed the validity of the diffusion approximation for the encoding current, which is illustrated for the 1st simulation in the inset of panel F of Fig. 3: the numerical estimation of the standard deviation of the spike count variance (performed through the Jack-Knife method for increasingly large time window size $\Delta_{\rm J}$) is, indeed, statistically consistent with its theoretical value $({\sigma_{\rm I}}^{2}+{\sigma_{\rm c}}^{2})^{1/2}$.